solve(d^2+dd'-6d'^2)z=cos(2x+y)


1
Expert's answer
2021-08-09T16:03:58-0400

The auxiliary equation is; "m^{2}+m-6=0\newline (m-2)(m+3)=0\newline m=2,\ -3"

C.F = f1(y+2x)+f2(y-3x)

P.I = "\frac{1}{D^{2}+DD^{I}-6{D^{I}}^{2}}cos(2x+y)"

=R.P"\frac{1}{(D-2D^{I})(D+3D^{I})}e^{i(2x+y)}"

=R.P"\frac{1}{(2i+3(i))}e^{i(2x+y)}"

=R.P"\frac{-i}{5}x(cos(2x+y)+i\ sin(2x+y))"

="\frac{-x}{5}cos(2x+y)"

z=f1(y+2x)+f2(y-3x)-"\frac{x}{5}cos(2x+y)"